OpenStudy (nuccioreggie):

What is the first step needed to solve 3 over 4 multiplied by x minus 3 equals negative 18? Add 3 to both sides Subtract 18 from both sides Multiply both sides by 3 Divide both sides by 4

pooja195 (pooja195):

\[\frac{ 3 }{ 4 } \times X-3=18\]

OpenStudy (nuccioreggie):

so B? @pooja195

pooja195 (pooja195):

theres something else you can do

OpenStudy (nuccioreggie):

o i think its A though

pooja195 (pooja195):

Yes! good job :)

OpenStudy (nuccioreggie):

ok thank you can you help me with a couple more please

pooja195 (pooja195):

sure :)

OpenStudy (nuccioreggie):

ok ill tag you